# LightRetriever: A LLM-based Text Retrieval Architecture with Extremely Faster Query Inference

### v4

@@ 1 Introduction
Given search queries and candidate documents, text retrieval employs dual-encoder architectures to encode representation vectors and search documents relevant to queries based on query-document similarities [63]. The encoded vectors could be condensed into low-dimensional vectors for dense retrieval or sparsified into high-dimensional vectors for sparse retrieval. Dense retrieval [30] compresses textual information into dense vectors with parameterized language models (LMs) and retrieves based on inner product similarity. Meanwhile, sparse retrieval [61] characterizes sparse term frequencies on a much broader vocabulary space, and then searches based on lexical overlap and term-based relevance metrics, such as direct impact search (Inner product) [3, 21] and BM25 [61]. Additionally, hybrid retrieval enhances retrieval abilities by interpolating their similarity scores [73] or ranks [15]. Text retrieval is broadly applied in various scenarios, such as web search [44], semantic textual similarity [8], fact checking [67], and retrieval augmented generation [40].
@@
Deploying retrieval systems requires different processing workflows for queries and documents. As is shown in Figure 1, as search candidates, documents are encoded with a document encoder $Enc_{d}$ and then indexed into search databases [28, 61]. Because documents are obtained before the search operation, they are pre-computable and distributable, which can be indexed offline and distributed anywhere [80]. On the contrary, queries are input and encoded in real-time, which are non-pre-computable and non-distributable. The query encoder $Enc_{q}$ needs to be served online. Special GPU accelerators are mandatory if the query encoder is a deeply parameterized LM. Most existing works of text retrieval [30, 11, 72] utilize symmetric dual-encoder architectures, which share the same LM for both the query and document encoding or have an equal number of parameters on both sides. Recent state-of-the-art (SOTA) retrievers [72, 5] employ deeply parameterized LLMs as dual-encoders $Enc_{q}$ and $Enc_{d}$ for enhanced retrieval capabilities and domain adaptation abilities. However, serving giant LLMs slows down query inference throughput and depletes online deployment resources [38]. Despite the rapid progress on improving retrieval quality, little prior work has investigated the efficiency challenges of LLM-based retrievers, necessitating special optimizations for query inference efficiency.
@@
Modern Transformer-based LMs empower text retrievers with rich contextual modeling through the attention mechanism [69], enabling deep interactions across contextual tokens. This capability explains the success of recent LLM-based dual-encoders [72, 5], but it also means that both queries and documents incur the same computational burden. In contrast, traditional statistical methods such as BM25 [61] rely purely on lexical overlap and token frequency, requiring negligible inference cost while still remaining competitive in many retrieval tasks [66]. This contrast highlights a practical tension: While document encoders can afford heavy pre-computation offline, query encoders must operate under strict online latency and resource constraints. Therefore, it is natural to ask: do queries truly need the same degree of deep contextual modeling as documents?
@@
To strike a balance between performances and efficiencies, we propose LightRetriever, a new hybrid retrieval architecture that explicitly breaks the symmetry between document and query encoders. Our key insight is that while documents benefit from the full modeling power of LLMs, queries do not require equally heavy processing to achieve strong retrieval performance. To this end, LightRetriever preserves a full-sized LLM encoder on the document side, but entirely removes the deep modeling on the query side. 1) For dense retrieval, we end-to-end train cacheable token embeddings from the LLM-based query encoder and aggregate them through simple averaging, reducing query inference to a single embedding lookup; 2) For sparse retrieval, we directly map token counts to sparse vectors without any encoding. Dense and sparse scores are summed [73] for the final hybrid retrieval. LightRetriever achieves retrieval quality comparable to full dual-encoder LLMs, but with orders-of-magnitude faster query inference.
@@
LightRetriever is super fast and lightweight for query inference. Compared to serving a full-sized LLM on an A800 GPU, our method achieves over a thousand times speedup for query encoding, and over 10x total throughput improvements. Experiments on large-scale retrieval benchmarks demonstrate that our method generalizes well across diverse LLMs (including Llama-1B, 3B, 8B, and Qwen-1.5B, 3B, 7B) and retrieval benchmarks (including 15 English tasks from BeIR [66] and 8 Chinese tasks from CMTEB Retrieval [77]), maintaining an average of 95% retrieval performance.
@@ 2 Algorithm
Figure 2: Contrastive training of full-sized symmetric dense retrieval. A full-sized dual-encoder is used to model both queries and documents. Task-specific instructions are added as common practices [65, 72] to promote better domain adaptation abilities.
@@ 2.1 LM-based text retrieval
which lowers the upper boundary of ranking similarity errors $\mathcal{E}$ [12]. The above objective can be effectively optimized by the pairwise/listwise contrastive loss [10], margin loss [62], etc. Following [30], listwise contrastive loss ($\ell^{CL}$) is used in our work,
@@
Similar to instructed fine-tuning (SFT) [56], modern LM-based retrievers [65, 77, 71], especially LLM retrievers [72, 5, 39], customize task specific instructions for better domain adaption abilities. Specifically, as shown in Figure 2, the task instruction is prepended to the queries before retrieval. Although [65, 77] use instructions on both query and passage sides, most LLM-based retrievers [72, 5, 39] do not use passage instructions to reuse the pre-built indexes across different tasks. In our work, dense retrieval follows the query instructions of Mistral-E5 [72] and does not involve passage instructions. The sparse retrieval is incompatible with instructions because it does not involve learnable query encoders. Detailed instructions are shown in the Appendix A.6.
@@ 2.2 Dense retrieval of LightRetriever
Given a task instruction $Inst$ and query tokens $T_{q}=\{t_{0},t_{1},...,t_{n-1}\}$, each query token $t_{i}$ ($i\in\{0,...,n-1\}$) are prepended with the instruction. Then, the corresponding query token vectors $v_{t_{i}}^{\text{den}}$ are obtained by encoding [^1] Specifically, EOS pooling is used in our work by following Mistral-E5 [72]. with the query encoder [^2] A customized causal mask is designed to avoid repeated prompt computations, detailed in Appendix A.7.,
@@
Query tokens do not deeply interact with each other. Thus, the query token vectors are cacheable. Given a certain vocabulary with size $V$, all token vectors $v_{t}^{den},t\in\{0,...,V-1\}$ are precomputed as in Equation 3 and stored in a lookup Embedding $E$.
@@
LLM-based encoders are no longer needed during online serving. Instead, only one Embedding layer, with a single weight matrix shaped $[V,H]$, is hosted in RAM. $H$ is the dimension of LLM’s hidden states, which is also the dimension of dense vectors. The dense query vector is obtained with a simple lookup-then-average operation,
@@
The online serving of our dense retriever is extremely lightweight, because such Embedding lookup is super efficient with or without GPU hardware.
@@ 2.3 Sparse retrieval of LightRetriever
where $t\in\{0,...,V-1\}$ represents valid token ids within LLM’s vocabulary. We obtain the sparse document vector $v_{d}^{\text{spr}}$ by first feeding the document $d$ into the pretrained language model (LM), resulting in the final-layer hidden states $h_{\text{last}}=LM(d)$. These hidden states are then projected into the vocabulary space to get logits $w$ via the model’s own language modeling head, parameterized by a projection matrix $P$. To induce sparsity and suppress dominant term frequencies, we apply a ReLU activation followed by a log-saturation function, and then perform a max-pooling operation of the logits $w$ over the sequence length dimension by following [21],
@@
$v_{d}^{\text{spr}}$ is not naturally sparse enough, because LM tends to generate dense distributions. Thus, the sparsify regulator is needed during the training phase. Given a batch of documents $d=\{d_{0},...,d_{i}\},i<\text{batch size }N$, the FLOPs regulator [57] is used for sparsification. This regulator first averages the logits across different documents within a certain batch, then computes the squared sum to get the final regulator.
@@
Table 1: Comparison of retrieval speed and effectiveness. We compare the time consumption of retrieving 65,536 Bing queries over 1 million passages from the MS MARCO dataset [53], evaluating both full symmetric retrievers and the query-lightweight LightRetriever. To assess the trade-off between efficiency and effectiveness, we also report nDCG@10 scores on English (BeIR) and Chinese (CMTEB-Retrieval) benchmarks. For retrieval efficiency, we report the consumed time of: query tokenization, model encoding, maximum search time of Faiss/Lucene, and total end-to-end retrieval time. The overall throughput is measured in Queries Per Second (QPS).
@@
LightRetriever-Llama8b 54.4 63.0 0.8209 0.0412 8.5010 9.3630 6999
Full-Qwen7b 56.6 70.1 1.4144 100.6716 8.5279 110.6140 592
Full-Qwen1.5b 54.0 65.9 1.4210 53.0858 8.5299 63.0370 1040
LightRetriever-Qwen7b 53.8 66.5 0.8300 0.0420 8.5150 9.3870 6982
3 Experiments
3.1 Settings
Training data
Our work finetunes on a large existing collection of 20 English and 3 Chinese datasets with 8.38M samples to ensure broad domain coverage and diversity. To ensure reproducibility, our work reuses the preprocessed fine-tuning data from tDRO [48] and Sentence Transformers Training Data [60]. English datasets include Amazon Review [54], Eli5 [20], FEVER [67], FiQA [49], GooQA [31], HotpotQA [81], MSMARCO [53], NFCorpus [7], NPR [45], NQ [37], PAQ [41], Quora Duplicates [26], S2ORC [43], Scifact [70], Specter [13], StackExchange Duplicates [17], E5 synthetic [4], Trivia [29], WikiHow [34], and Yahoo Answers [83]. Chinese datasets include cMedQA2 [82], DuReader [58], and T2Ranking [78].
Benchmarks
BeIR [66] and CMTEB Retrieval [77] benchmarks are used in the evaluation. BeIR is a massive English retrieval benchmark collection with 15 evaluation sets. These heterogeneous datasets cover different domains, tasks, and retrieval granularities. CMTEB Retrieval is a massive Chinese retrieval benchmark collection with 8 sets, also covering a variety of domains. nDCG@10 is reported as the main metric by following [66, 77], accessing both recall and ranking abilities. Recalls (R@{20, 50, 100}) are also reported to assess retrieval ability at large windows.
Training hyper-parameters
Our experiments are conducted based on the implementation of tDRO [48]. All experiments are performed with a batch size of 128, 7 hard negatives, a max sequence length of 512, a contrastive temperature $\tau$ of 0.02, and 12k total steps. Following SPLADE [21], the coefficient of the FLOPs regulator is 0.001, and the regulator is quadratically increased to the maximum at the first 4k steps, which helps reduce the side effects of the regulator at initial steps [57]. LoRA [25] is used to save training GPU memory, where LoRA r is 16, alpha is 32, and dropout is 0.1. Different LLMs with different sizes are tested as the backbone encoders to verify the generalization abilities, including Llama-3.2-1B, 3B, Llama-3.1-8B [68], and Qwen-2.5-1.5B, 3B, 7B [2]. The hybrid similarity scores are linearly summed from normalized dense and sparse scores [73]. All trainings are conducted on 8 NVIDIA H800 or A800 GPUs.
Speed comparisons
To evaluate the online speedup of LightRetriever, we compare the online time consumptions for retrieving 65536 Bing queries over 1M passages from MSMARCO[53]. All speed tests are conducted on a single A800 GPU, with a batch size of 256, a dimension of 1k for Faiss exact search, and 64 threads for Anserini (Lucene) sparse search. Faiss and Anserini search in parallel, where Faiss searches more slowly in our cases.
@@ 3.1 Settings
Performance comparisons
@@
Notably, Static Embedding performs retrieval using a single shared embedding layer for both queries and documents, offering low retrieval performance but highly efficient baseline. BM25, as a classical term-frequency-based method, remains a widely adopted sparse retriever. The combination of them thus provides a reference for efficient retrieval on both sides. In contrast, BGE-m3 is a Transformer-based retriever trained with multi-stage pretraining and fine-tuning, and represents state-of-the-art performance in hybrid retrieval.
Except for Static Embedding and BM25, all compared retrievers rely on deep Transformer encoders on both query and document sides. LightRetriever is, to our knowledge, the first method to address the online efficiency bottleneck on the query side of LLM-based retrievers, striking a balance between retrieval quality and inference efficiency.
3.2 Main results
Encoding speeds
In contrast, LightRetriever reduces encoding time to below 50 ms (0.0412–0.0420s), achieving over a 2500× speedup in the encoding phase. This drastic reduction is made possible by replacing full forward passes of large Transformer models with a simple embedding lookup. And the overall throughput achieves over 10x QPS speedup. Notably, this lightweight encoding does not significantly compromise retrieval performance. LightRetriever maintains competitive nDCG@10 scores on both English (BeIR) and Chinese (CMTEB-Retrieval) benchmarks.
These results prove the effectiveness of LightRetriever’s decoupled query modeling paradigm, which drastically improves online query encoding efficiency while preserving good retrieval quality, making it suitable for latency-critical applications.
Performance comparisons
Table 2 compares retrieval effectiveness on English (BeIR) and Chinese (CMTEB-Retrieval) benchmarks. The main focus of this comparison is to evaluate the performance degradation introduced by LightRetriever’s lightweight query encoding. Overall, by leveraging deep Transformer encoders on both query and document sides, full symmetric retrievers with deep LLM encoders achieve the best scores (e.g., Llama3.1-8b on BeIR with 56.8 nDCG@10, Qwen2.5-7b on CMTEB-R with 70.1 nDCG@10), but at significant computational cost.
LightRetriever achieves promising effectiveness while drastically reducing query-side computational burden. Across different backbones, it incurs only modest degradations (typically around 5% with 1–5 absolute points). For example, LightRetriever-Qwen2.5-7b achieves 53.8 nDCG@10 on BeIR and 66.5 nDCG@10 on CMTEB-R, only 2.8 and 3.6 lower than its full counterpart. Detailed task-level results are shown in Appendix Tables 911. Notably, LightRetriever retains strong performance on large-scale open-domain QA, fact-checking, and web search benchmarks such as FEVER [67], SciFact [70], TRECCOVID [16], and MSMARCO [53]. On smaller, domain-specific datasets (e.g., FiQA [49] and CmedQA [82]), drops can be moderately larger (around 5 points) but remain tolerable, suggesting good generalization with room for future task-specific adaptation.
[…]

### v5

@@ 1 Introduction
Given search queries and candidate documents, text retrieval employs dual-encoder architectures to encode representation vectors and search documents relevant to queries based on query-document similarities (Salton et al., 1975). The encoded vectors could be condensed into low-dimensional vectors for dense retrieval or sparsified into high-dimensional vectors for sparse retrieval. Dense retrieval (Karpukhin et al., 2020) compresses textual information into dense vectors with parameterized language models (LMs) and retrieves based on inner product similarity. Meanwhile, sparse retrieval (Robertson et al., 1994) characterizes sparse term frequencies on a much broader vocabulary space, and then searches based on lexical overlap and term-based relevance metrics, such as direct TF-based search (Inner product) (Bai et al., 2020; Formal et al., 2021) and BM25 (Robertson et al., 1994). Additionally, hybrid retrieval enhances retrieval abilities by interpolating their similarity scores (Wang et al., 2021) or ranks (Cormack et al., 2009). Text retrieval is broadly applied in various scenarios, such as web search (Lu et al., 2022), semantic textual similarity (Bowman et al., 2015), fact checking (Thorne et al., 2018), and retrieval augmented generation (Lewis et al., 2020).
@@
Deploying retrieval systems requires different processing workflows for queries and documents. As shown in Figure 1, as search candidates, documents are encoded with a document encoder $Enc_{d}$ and then indexed into search databases (Johnson et al., 2021; Robertson et al., 1994). Because documents are obtained before the search operation, they are pre-computable and distributable, which can be indexed offline and distributed anywhere (Yang et al., 2018a). On the contrary, queries are input and encoded in real-time, which are non-pre-computable and non-distributable. The query encoder $Enc_{q}$ needs to be served online. Special GPU accelerators are mandatory if the query encoder is a deeply parameterized LM. Recent state-of-the-art (SOTA) retrievers (Wang et al., 2024a; BehnamGhader et al., 2024) employ deeply parameterized LLMs as symmetric dual-encoders $Enc_{q}$ and $Enc_{d}$ for enhanced retrieval capabilities. However, serving giant LLMs slows down query inference throughput and depletes online deployment resources (Kwon et al., 2023). Despite the rapid progress on improving retrieval quality, little prior work has investigated the efficiency challenges of LLM-based retrievers, necessitating special optimizations for query inference efficiency.
@@
Modern Transformer-based LMs empower text retrievers with rich contextual modeling through the attention mechanism (Vaswani et al., 2017), enabling deep interactions across contextual tokens. This capability explains the success of recent LLM-based dual-encoders (Wang et al., 2024a; BehnamGhader et al., 2024), but it also means that both queries and documents incur the same computational burden. In contrast, traditional statistical methods such as BM25 (Robertson et al., 1994) rely purely on lexical overlap and token frequency, requiring negligible inference cost while still remaining competitive in many retrieval tasks (Thakur et al., 2021). This contrast highlights a practical tension: While document encoders can afford heavy pre-computation offline, query encoders must operate under strict online latency and resource constraints. Therefore, it is natural to ask: do queries truly need the same degree of deep contextual modeling as documents?
@@
To strike a balance between performances and efficiencies, we propose LightRetriever, a new hybrid retrieval architecture that explicitly breaks the symmetry between document and query encoders. Our key insight is that while documents benefit from the full modeling power of LLMs, queries do not require equally heavy processing to achieve strong retrieval performance. To this end, LightRetriever preserves a full-sized LLM encoder on the document side, but entirely removes the deep modeling on the query side. 1) For dense retrieval, we end-to-end train cacheable token embeddings from the LLM-based query encoder and aggregate them through simple averaging, reducing query inference to a single embedding lookup; 2) For sparse retrieval, we directly map token counts to sparse vectors without any encoding. Dense and sparse scores are linearly interpolated (Wang et al., 2021) for the final hybrid retrieval.
@@
LightRetriever achieves retrieval quality mostly comparable to full dual-encoder LLMs, but with orders-of-magnitude faster query inference. Compared to serving a full-sized LLM on an A800 GPU, our method achieves over a thousand times speedup for query encoding, and over 10x total throughput improvements. Importantly, our experiments show that adopting a lightweight query encoder does not mean queries are deprived of deep semantic understanding. Our work tries to transfer the major modeling cost from the query side to the document side, while still preserving the benefits of LLM-powered relevance-based similarity search with query-document interactions. Experiments on large-scale retrieval benchmarks demonstrate that our method generalizes well across diverse LLMs (including Llama-1B, 3B, 8B, and Qwen-1.5B, 3B, 7B) and retrieval benchmarks (including 15 English tasks from BeIR (Thakur et al., 2021) and 8 Chinese tasks from CMTEB Retrieval (Xiao et al., 2024)), maintaining an average of 95% retrieval performance.
@@ 2 Algorithm
Figure 2: Contrastive training of full-sized symmetric dense retrieval. A full-sized dual-encoder is used to model both queries and documents. Task-specific instructions are added as common practices (Su et al., 2023; Wang et al., 2024a) to promote better domain adaptation abilities.
@@ 2.1 LM-based text retrieval
which lowers the upper boundary of ranking similarity errors $\mathcal{E}$ (Chen et al., 2009). The above objective can be effectively optimized by contrastive loss (Cao et al., 2007), margin loss (Rosset et al., 2003), etc. Following (Karpukhin et al., 2020), listwise contrastive loss ($\ell^{CL}$) is used in our work,
@@
Similar to instructed fine-tuning (SFT) (Ouyang et al., 2022), modern LM-based retrievers (Su et al., 2023; Xiao et al., 2024; Wang et al., 2024b), especially LLM retrievers (Wang et al., 2024a; BehnamGhader et al., 2024; Lee et al., 2024), customize task specific instructions for better domain adaption abilities. Specifically, as shown in Figure 2, the task instruction is prepended to the queries before retrieval. Although (Su et al., 2023; Xiao et al., 2024) use instructions on both query and passage sides, most LLM-based retrievers (Wang et al., 2024a; BehnamGhader et al., 2024; Lee et al., 2024) do not use passage instructions to reuse the pre-built indexes across different tasks. In our work, dense retrieval follows the query instructions of Mistral-E5 (Wang et al., 2024a) and does not involve passage instructions. The sparse retrieval is incompatible with instructions because it does not involve learnable query encoders. Detailed instructions are shown in the Appendix A.7.
@@ 2.2 Dense retrieval of LightRetriever
Given a task instruction $Inst$ and query tokens $T_{q}=\{t_{0},t_{1},...,t_{n-1}\}$, each query token $t_{i}$ ($i\in\{0,...,n-1\}$) are prepended with the instruction. Then, the corresponding query token vectors $v_{t_{i}}^{\text{den}}$ are obtained via last token pooling with the query encoder [^1] A customized causal mask is designed to avoid repeated prompt computations, detailed in Appendix A.8.. Then, the dense query vector $v_{q}^{\text{den}}$ of the whole query text is aggregated by averaging token vectors,
@@
Query tokens do not deeply interact with each other. Thus, the query token vectors are cacheable. Given a certain vocabulary with size $V$, all token vectors $v_{t}^{den},t\in\{0,...,V-1\}$ are precomputed as in Equation 3 and stored in a lookup Embedding $E$. Typically, caching such a lookup Embedding on 8 H800 with a Llama-8b model consumes less than 20s, which is a one-go and cheap operation. In Appendix A.4, LightRetriever also shows good ability to balance the serving Embedding size via simple dimension truncations (Kusupati et al., 2022).
@@
LLM-based encoders are no longer needed during online serving. Instead, only one Embedding layer, with a single weight matrix shaped $[V,H]$, is hosted in RAM. $H$ is the dimension of LLM’s hidden states, which is also the dimension of dense vectors. The dense query vector is obtained with a simple lookup-then-average operation, which is super efficient with or without GPU acceleration.
@@ 2.3 Sparse retrieval of LightRetriever
where $t\in\{0,...,V-1\}$ represents valid token ids within LLM’s vocabulary. We obtain the sparse document vector $v_{d}^{\text{spr}}$ by first feeding the document $d$ into the pretrained language model (LM), resulting in the final-layer hidden states $h_{\text{last}}=LM(d)$. These hidden states are then projected into the vocabulary space to get logits $w$ via the model’s own language modeling head, parameterized by a projection matrix $P$. To induce sparsity and suppress dominant term frequencies, we apply a ReLU activation followed by a log-saturation function, and then perform a max-pooling operation of the logits $w$ over the sequence length dimension by following (Formal et al., 2021),
@@
$v_{d}^{\text{spr}}$ is not naturally sparse enough, because LM tends to generate dense distributions. Thus, the sparsify regulator is needed during the training phase. Given a batch of documents $d=\{d_{0},...,d_{i}\},i<\text{batch size }N$, the FLOPs regulator (Paria et al., 2020) is used for sparsification. This regulator first averages the logits across different documents within a certain batch, then computes the squared sum to get the final regulator.
@@
3 Experiments
3.1 Settings
Training data
Our work finetunes on a large existing collection of 20 English and 3 Chinese datasets with 8.38M samples to ensure broad domain coverage and diversity. To ensure reproducibility, our work reuses the preprocessed fine-tuning data from tDRO (Ma et al., 2025) and Sentence Transformers Training Data. English datasets include Amazon Review (Ni et al., 2019), Eli5 (Fan et al., 2019), FEVER (Thorne et al., 2018), FiQA (Maia et al., 2018), GooAQ (Khashabi et al., 2021), HotpotQA (Yang et al., 2018b), MSMARCO (Nguyen et al., 2016), NFCorpus (Boteva et al., 2016), NPR (Lucas et al., 2023), NQ (Kwiatkowski et al., 2019), PAQ (Lewis et al., 2021), Quora Duplicates (Iyer et al., 2017), S2ORC (Lo et al., 2020), Scifact (Wadden et al., 2020), Specter (Cohan et al., 2020), StackExchange Duplicates (da Silva et al., 2018), E5 synthetic (BeastyZ, 2024), Trivia (Joshi et al., 2017), WikiHow (Koupaee and Wang, 2018), and Yahoo Answers (Zhang et al., 2015). Chinese datasets include cMedQA2 (Zhang et al., 2018), DuReader (Qiu et al., 2022), and T2Ranking (Xie et al., 2023).
Benchmarks
BeIR (Thakur et al., 2021) and CMTEB Retrieval (Xiao et al., 2024) benchmarks are used in the evaluation. BeIR is a massive English retrieval benchmark collection with 15 evaluation sets. These heterogeneous datasets cover different domains, tasks, and retrieval granularities. CMTEB Retrieval is a massive Chinese retrieval benchmark collection with 8 sets, also covering a variety of domains. nDCG@10 is reported as the main metric by following (Thakur et al., 2021; Xiao et al., 2024), accessing both recall and ranking abilities. Recalls (R@{20, 50, 100}) are also reported to assess retrieval ability at large windows.
Training hyper-parameters
Our experiments are conducted based on the implementation of tDRO (Ma et al., 2025). All experiments are performed with a batch size of 128, 7 hard negatives, a max sequence length of 512, a contrastive temperature $\tau$ of 0.02, and 12k total steps. Following SPLADE (Formal et al., 2021), the coefficient of the FLOPs regulator is 0.001, and the regulator is quadratically increased to the maximum at the first 4k steps, which helps reduce the side effects of the regulator at initial steps (Paria et al., 2020). LoRA (Hu et al., 2022) is used to save training GPU memory, where LoRA r is 16, alpha is 32, and dropout is 0.1. Different LLMs with different sizes are tested as the backbone encoders to verify the generalization abilities, including Llama-3.2-1B, 3B, Llama-3.1-8B (Touvron et al., 2023), and Qwen-2.5-1.5B, 3B, 7B (Bai et al., 2023). The hybrid similarity scores are linearly summed from normalized dense and sparse scores (Wang et al., 2021). All trainings are conducted on 8 NVIDIA H800 or A800 GPUs. Specifically, training a LightRetriver-Llama8b on 8 H800 consumes about 15 hours.
Speed comparisons
To evaluate the online speedup of LightRetriever, we compare the online time consumptions for retrieving 65536 Bing queries over 1M passages from MSMARCO (Nguyen et al., 2016). All speed tests are conducted on a single A800 GPU, with a batch size of 256, a dimension of 1k for Faiss exact search, and 64 threads for Anserini (Lucene) sparse search. Faiss and Anserini search in parallel, where Faiss searches more slowly in our cases.
Performance comparisons
[…]
Notably, Static Embedding performs retrieval using a single shared embedding layer for both queries and documents, offering low retrieval performance but a highly efficient baseline. BM25, as a classical term-frequency-based method, remains a widely adopted sparse retriever. The combination of them thus provides a reference for efficient retrieval on both sides. In contrast, BGE-m3 is a Transformer-based retriever trained with multi-stage pretraining and fine-tuning, and represents state-of-the-art performance in hybrid retrieval. USTAD (Kim et al., 2024) reports a small 6-layer DistilBERT retriever, offering a SOTA distillation baseline for small language models (SLM). In Appendix A.5, applications of similar distillation techniques are also discussed in our works.
Except for Static Embedding and BM25, all compared retrievers, especially LLM retrievers, rely on deep Transformer encoders on both query and document sides. LightRetriever is, to our knowledge, the first method to address the online efficiency bottleneck on the query side of LLM-based retrievers, striking a balance between retrieval quality and inference efficiency.
3.2 Main results
Encoding speeds
In contrast, LightRetriever reduces encoding time to below 50 ms (0.0412–0.0420s), reaching a maximal 2500× speedup of full-sized Llama-8b in the encoding phase. This drastic reduction is made possible by replacing full forward passes of large Transformer models with a simple embedding lookup. And the overall throughput achieves over 10x QPS speedup. Notably, this lightweight encoding does not significantly compromise retrieval performance. LightRetriever maintains competitive nDCG@10 scores on both English (BeIR) and Chinese (CMTEB-Retrieval) benchmarks.
A potent baseline to improve LLM query encoder efficiency is layer reduction. We evaluated by training and serving only the first Transformer layer of the Llama3-8b model on the query side under the same settings, meanwhile retaining full-sized model on the document side. While this reduces model size during training, it underperforms the LightRetriever (–4.3 on BeIR, –8.6 on CMTEB-R) and introduces inference-time overhead of 2.34s. This proves the need for full-sized query modeling during the training phase.
The above results prove the effectiveness of our extreme decoupled modeling paradigm, which drastically improves online query encoding efficiency while preserving retrieval quality, making it suitable for latency-critical applications.
[…]
@@ 3.2 Main results
[…]
LightRetriever-Llama8b 54.4 63.0 0.8209 0.0412* 8.5010 9.3630 6999*
@@
Overall performance comparisons
Table 2 compares retrieval effectiveness on English (BeIR) and Chinese (CMTEB-Retrieval) benchmarks. The main focus of this comparison is to evaluate the performance degradation introduced by LightRetriever’s lightweight query encoding. Overall, by leveraging deep Transformer encoders on both query and document sides, full symmetric retrievers with deep LLM encoders achieve the best scores (e.g., Llama3.1-8b on BeIR with 56.8 nDCG@10, Qwen2.5-7b on CMTEB-R with 70.1 nDCG@10), but at a significant computational cost.
LightRetriever achieves promising effectiveness while drastically reducing query-side computational burden. Across different backbones, it incurs only modest degradations (typically around 5% with 1–5 absolute points). For example, LightRetriever-Qwen2.5-7b achieves 53.8 nDCG@10 on BeIR and 66.5 nDCG@10 on CMTEB-R, only 2.8 and 3.6 lower than its full counterpart.
@@
Category-wise performance comparison
To provide a more fine-grained analysis of our query-lightweight proposal across different retrieval scenarios, we further categorize BeIR datasets into 8 task groups by following the original paper (Thakur et al., 2021) and summarize the results in Table 3. Overall, LightRetriever shows strong robustness on general relevance-based retrieval tasks, including Bio-Medical IR, Argument Retrieval, Fact Checking, Duplicate Question Retrieval, Passage Retrieval, and General QA, where it consistently preserves over 93% of the full symmetric retriever’s effectiveness, and even surpasses it in some cases (e.g., TREC-COVID, ArguAna, and Touche2020). This indicates that lightweight query encoding is sufficient to capture the semantic signals in typical relevance-based ad-hoc and open-domain retrieval.
In contrast, moderately larger performance degradations are observed on more challenging tasks, such as Domain-specific QA (HotpotQA, FiQA), Entity Retrieval (DBPedia), and Citation Prediction (SCIDOCS), where the retention drops to around 87-89%. Additional benchmarks on CoIR (Li et al., 2025) and BRIGHT (Su et al., 2025) in §A.10 also show inferior performance of query-lightweight design on code and reasoning-intensive retrieval. These tasks typically require higher reasoning and query-understanding ability, which is not suitable for an extreme lightweight query encoder. Nevertheless, LightRetriever still maintains competitive absolute performance, suggesting promising generalization with room for further task-specific adaptation.
Table 3: Category-wise performance comparison on BEIR. We report the average nDCG@10 of the full symmetric retriever (Llama-3.1-8B), LightRetriever (Llama-3.1-8B), and the performance retention of LightRetriever (LightRetriever/Full-Symmetric).
Task Datasets Full Light Retention
Bio-Medical IR TREC-COVID 62.0 71.5 115.3%
Argument Retrieval ArguAna, Touche2020 39.7 42.5 107.2%
Fact Checking FEVER, ClimateFEVER, SciFact 70.8 67.6 95.5%
Duplicate Question Retrieval CQADupstack, Quora 69.6 65.3 93.8%
Passage Retrieval / General QA MSMARCO, NFCorpus, NQ 50.1 47.0 93.8%
Domain-specific QA HotpotQA, FiQA2018 68.4 61.1 89.3%
Entity Retrieval DBPedia 48.2 42.0 87.1%
Citation Prediction SCIDOCS 23.8 20.7 87.0%